If you have 2.5 mols of magnesium how many moles of magnesium nitride can be produced


1
Expert's answer
2021-01-11T03:59:45-0500

3Mg+N2=Mg3N2.n(Mg3N2)=1/3n(Mg)=1/3∗2.5=0.83mol